an airplane can average 100mph. if a trip take 2 hours one way and return takes 1 hour and 15 minutes, find the wind speed assuming it is constant.

Answer by stanbon(75887)

an airplane can average 100mph.
if a trip take 2 hours one way and
return takes 1 hour and 15 minutes,
find the wind speed assuming it is constant.
-----
Going DATA:
time = 2 hrs ; rate = 100-w ; distance = 2(100-w) miles
---
Return DATA:
time = (5/4) hrs ; rate = 100+w ; distance = (5/4)(100+w) miles
----
Equation:
distance = distance
200-2w = (5/4)100 + (5/4)w
---
800 - 8w = 500 + 5w
300 = 13w
----
w = 23.08 mph (wind speed)
=============================
Cheers,
Stan H.
============